# XL-DURel: Finetuning Sentence Transformers for Ordinal Word-in-Context Classification

### v1

@@ 5 Evaluation
Following the the CoMeDi shared task, we use ordinal Krippendorff’s $\alpha$ (Krippendorff, 2018) as evaluation measure for ordinal classification. It penalizes stronger deviations from the gold label more heavily. It has the additional advantage of controlling for expected disagreement and has been demonstrated to be superior to other measures such as Mean Absolute Error for ordinal classification (Sakai, 2021). We also use Spearman’s rank correlation coefficient ($\rho$) between continuous cosine similarities and gold ordinal labels to assess the rank alignment of model predictions. This enables us to evaluate performance without inducing thresholds. We further apply the nominal version of Krippendorff and the Spearman correlation for binary label evaluation. During training cosine similarities are compared to binary labels with the Average Precision.
@@ 6 Experiments
We now test which loss functions (see Section 4.2) and data combinations (see Section 3) improve performance on the ordinal CoMeDi test data over the baselines described in Section 4.5. We additionally report performance on binary WiC data to understand whether optimization for ordinal data hurts the binary task. As finetuning is computationally expensive, we perform one run for each model. [^10] We selectively re-ran models and observed variation of average performances between ±.$\pm.01$$03.$
Models are evaluated with Krippendorff’s $\alpha$ as described in Section 5 based on their cosine predictions binned to ordinal labels (see Section 4.4). [^11] We also experimented with using the angle difference on test for AnglE Loss, but did not outperform the cosine similarity.  [^12] In Appendix C, we give an additional performance evaluation on the raw cosine predictions with Spearman rank correlation. All experiments follow the training setup described in Section 4.3.
@@
Results are shown in Figure 2. With a performance of .67, AnglE Loss achieves the best result on CoMeDi test data (‘CoMeDi-Avg’) when using both WiC and CoMeDi training data (AnglE-WiC+CoMeDi). With performances of .64 and .62, it is followed by the CoSENT Loss (CSENT-WiC+CoMeDi) and the Cosine Loss (COS-WiC+CoMeDi). These models clearly outperform the published version of XL-LEXEME (XL-LXM), its retrained model version (CL-WiC) and the retrained version with additional ordinal data (CL-WiC+CoMeDi) with performances of .58, .56 and .57 respectively. The top model AnglE-WiC+CoMeDi outperforms the latter ones on all languages and by a large margin on average. Consequently, it clearly outperforms the published XL-LEXEME result from CoMeDi the shared task (XL-LXM-CoMeDi) .67 vs. .58. It further slightly outperforms the shared task winning model DeepMistake (.67 vs. .66). This is also the case if we exclude Spanish (‘CoMeDi-Avg-W/O-ES’), which was reported additionally by the task organizers because DeepMistake was trained on part of the test data for Spanish. Notably, we reach this performance by optimizing one unified model while the DeepMistake result was obtained by optimizing multiple models tailored to specific languages. [^13] We provide the finetuned AnglE-WiC+CoMeDi model at https://huggingface.co/sachinn1/xl-durel.
On binary WiC, performance remains more stable across models ranging between .62 and .68 (‘WiC-Avg’). Top performance is reached by the CoSENT Loss model relying on both binary and ordinal training data (CSENT-WiC+CoMeDi), closely followed by the AnglE Loss relying on the same data (AnglE-WiC+CoMeDi) and XL-LEXEME with .67 respectively.
@@
We observe that adding ordinal data improves performance on the ordinal task across loss functions. Compare for example performances of AnglE-WiC vs. AnglE-WiC+CoMeDi or CSENT-WiC vs. CSENT-WiC+CoMeDi. There is a clear average performance improvement on the ordinal task (‘CoMeDi-Avg’). Similarly for performance on the binary task (‘WiC-Avg’), but only for the ranking losses AnglE and CoSENT. However, for the Contrastive and the Cosine Similarity Loss the performance drop is negligible. These results suggest that binary WiC can be treated as a special case of ordinal WiC and that optimizing models for the general ordinal task improves performance on the more specific binary task.
More generally, Figure 2 shows that finetuning has a strong effect on performance. For this, compare all models against SBERT. Further, we find that training purely on CoMeDi yields competitive performance across the ordinal and the binary task (AnglE-CoMeDi). Binarizing ordinal training data and using Contrastive Loss (CL-CoMeDi) yields worse performance overall, even on the binary task. Training the top loss model purely on WiC (AnglE-WiC) underperforms on the ordinal task, but yields competitive performance on the binary task. Training on binary labels with the Contrastive Loss (CL-WiC) leads to rather worse performance than using AnglE Loss (AnglE-WiC). When training on WiC and CoMeDi labels, using AnglE Loss is clearly favorable (AnglE-WiC+CoMeDi vs. CL-WiC+CoMeDi) for both tasks. When using the Contrastive Loss, additional CoMeDi labels slightly help for the ordinal task, but not for the binary task (CL-WiC+CoMeDi vs. CL-WiC). However, when using the AnglE Loss, additional CoMeDi labels help clearly for both tasks (AnglE-WiC+CoMeDi vs. AnglE-WiC).
Our results are in line with the motivations given for the loss functions in Section 4.2: AnglE and CoSENT Loss are explicitly optimizing a ranking objective, which exploits the inherent ordering of ordinal labels. Further, the AnglE Loss improves optimization over the CoSENT Loss, presumably because it avoids killed gradients occurring with the cosine similarity.
@@ 7 Conclusion
We compared several loss functions for classification, regression and ranking tasks to finetune OGWiC models. Our top model outperformed previous models on ordinal and binary WiC data with a ranking objective based on angular distance in complex space. Overall, we find that using the AnglE Loss can be recommended, both for the ordinal and the binary WiC task. Similarly, mixing ordinal and binary training data is beneficial for both tasks. These results suggest that binary WiC can be treated as a special case of ordinal WiC and that optimizing models for the general ordinal task improves performance on the more specific binary task. In the future, WiC task setups should try to unify these approaches in order to make use of the full power of WiC training signals from multiple types of data. Further, we should try to optimize models more directly for ordinal classification instead of ranking (Vargas et al., 2020).
@@
We tested our hypotheses using particular data, base models and training architectures. In future research, these should be varied to test whether they have an influence on effects. Specifically, it should be tested whether the AnglE Loss turns out to be beneficial for the ordinal and binary task with additional test data. It is also unclear why the AnglE Loss performs better with the cosine similarity than with the angle difference in complex space although the latter is optimized during training.
@@ Appendix B CoMeDi Data
Figure 3: Model evaluation with Spearman on binary and ordinal test data. CL = Contrastive Loss, COS = Cosine Similarity Loss, CSENT = CoSENT Loss, AnglE = AnglE Loss, XL-LXM = XL-LEXEME.
Appendix C Spearman Results
Table 3 shows model performances measured with Spearman correlation.

### v2

@@ 5 Evaluation
Following the CoMeDi shared task, we use ordinal Krippendorff’s $\alpha$ (Krippendorff, 2018) as evaluation measure for ordinal classification. It penalizes stronger deviations from the gold label more heavily. It has the additional advantage of controlling for expected disagreement and has been demonstrated to be superior to other measures such as Mean Absolute Error for ordinal classification (Sakai, 2021). We also use Spearman’s rank correlation coefficient ($\rho$) between continuous similarities and gold ordinal labels to assess the rank alignment of model predictions. This enables us to evaluate performance without inducing thresholds. We further apply the nominal version of Krippendorff and the Spearman correlation for binary label evaluation. During training, similarities are compared to ordinal labels with Spearman correlation and to binary labels with the Average Precision.
@@ 6 Experiments
We now test which loss functions (see Section 4.2) and data combinations (see Section 3) improve performance on the ordinal CoMeDi test data over the baselines described in Section 4.5. We additionally report performance on binary WiC data to understand whether optimization for ordinal data hurts the binary task. As finetuning is computationally expensive, we perform one run for each model. [^11] We selectively re-ran models and observed variation of average performances between $\pm.01$.03.03. Models are evaluated with Krippendorff’s $\alpha$ as described in Section 5 based on their cosine predictions binned to ordinal labels (see Section 4.4). [^12] We also experimented with using the angle difference on test for AnglE Loss, but did not outperform the cosine similarity. All experiments follow the training setup described in Section 4.3. Results are shown in Figure 2.131313In Appendix D, we give an additional performance evaluation on the raw cosine predictions with Spearman rank correlation.
Loss function
@@
First of all, we see that finetuning has a strong effect on performance. For this, compare all models against SBERT, which is the only non-finetuned model. With a performance of .67, AnglE Loss achieves the best result on CoMeDi test data (‘CoMeDi-Avg’) when using both WiC and CoMeDi training data (AnglE-WiC+CoMeDi). With performances of .64 and .62 respectively, it is followed by the CoSENT Loss (CSENT-WiC+CoMeDi) and the Cosine Similarity Loss (COS-WiC+CoMeDi). All these models clearly outperform the published version of XL-LEXEME (XL-LXM), its retrained model version (CL-WiC) and the retrained version with additional ordinal data (CL-WiC+CoMeDi), which have performances of .58, .56 and .57, respectively. The top model AnglE-WiC+CoMeDi outperforms the latter ones on all languages and on average by a large margin. The same holds for the published XL-LEXEME result from the CoMeDi shared task (XL-LXM-CoMeDi), which is outperformed with .67 vs. .58. It further slightly outperforms the shared task winning model DeepMistake (.67 vs. .66). This is also the case if we exclude Spanish (‘CoMeDi-Avg-W/O-ES’), which was reported additionally by the task organizers because DeepMistake was trained on part of the test data for Spanish. Notably, we reach this performance by optimizing one unified model while the DeepMistake result was obtained by optimizing multiple models tailored to specific languages. [^14] We provide the finetuned AnglE-WiC+CoMeDi model under the name “XL-DURel” at https://huggingface.co/sachinn1/xl-durel. Find the code for reproducing our results at https://github.com/sachinn12/XL-DURel. On binary WiC (‘WiC-Avg’), top performance is reached by the CoSENT Loss model relying on both binary and ordinal training data (CSENT-WiC+CoMeDi) with .68, closely followed by the AnglE Loss relying on the same data (AnglE-WiC+CoMeDi) and XL-LEXEME with .67, respectively. Hence, on the ordinal and the binary task ranking losses show top performance, where the advantage to the classification loss is more pronounced for the ordinal task. While the regression loss is not competitive for either task, it shows a clear advantage over the classification loss on the ordinal task. These results are in line with the motivations given for the loss functions in Section 4.2: AnglE and CoSENT Loss are explicitly optimizing a ranking objective, which exploits the inherent ordering of ordinal labels. Further, the AnglE Loss improves optimization over the CoSENT Loss and other cosine-based losses, presumably because it avoids killed gradients occurring with the cosine similarity.
Training data
@@
Note that training on purely ordinal data yields good baseline performance across tasks, especially with the AnglE Loss (AnglE-CoMeDi). For binary data, this is also the case, but there is a larger performance difference to top models on the ordinal task (e.g. AnglE-WiC). Moreover, we clearly observe that combining ordinal and binary data improves performance on the ordinal task across all loss functions. Compare for example performances of AnglE-WiC/CoMeDi vs. AnglE-WiC+CoMeDi or CSENT-WiC/CoMeDi vs. CSENT-WiC+CoMeDi. There is a clear average performance improvement on the ordinal task. Similarly for performance on the binary task, but only for the ranking losses AnglE and CoSENT. However, for the Contrastive and the Cosine Similarity Loss the performance drop is negligible.
Ordinality
Regardless of training data, the ordinal training signal turns out to be beneficial for both tasks. In order to see this, compare e.g. CL-CoMeDi vs. AnglE-CoMeDi. The former binarizes the data while the latter keeps the ordinal information. The performance difference is striking with .58 vs. .62 for the ordinal task and, interestingly, .57 vs. .63 for the binary task. This indicates that fine-grained semantic proximity information helps the model to better learn binary meaning distinctions, which is also supported by the fact that the best purely ordinal model approaches the performance of the best purely binary model on the binary task (AnglE-CoMeDi vs. CL-WiC) despite being trained on much less and out-of-distribution data.
@@ 7 Conclusion
We compared several loss functions for classification, regression and ranking to finetune OGWiC models. Our top model outperformed previous models on ordinal and binary WiC data with a ranking objective based on angular distance in complex space. Overall, we found that using the AnglE Loss can be recommended, both for the ordinal and the binary WiC task. Similarly, mixing ordinal and binary training data turned out to be beneficial for both tasks. These results suggest that binary WiC can be treated as a special case of ordinal WiC and that optimizing models for the general ordinal task improves performance on the more specific binary task. In the future, WiC task setups should try to unify these approaches in order to make use of the full power of WiC training signals from multiple types of data. Further, we should try to optimize models more directly for ordinal classification instead of ranking. Currently, our model first predicts a dense similarity which we then discretize in an independent step to ordinal labels through thresholds. However, there are also loss functions directly optimizing for ordinal labels, like Cumulative Link models (Vargas et al., 2020). We would like to test such models for OGWiC. Moreover, similar to Loke et al. (2025), we would like to test Large Language Models such as Llama (Touvron et al., 2023). These profit from massive amounts of parameters and training data and can be directly instructed to predict an ordinal number. Note, however, that our current approach has certain advantages over this: It is theoretically motivated by employing direct ranking optimization. Also, it is small and efficient making the model applicable to large amounts of data.
@@
We tested our hypotheses using particular data, base models and training architectures. In future research, these should be varied to test whether they have an influence on effects. Specifically, it should be tested whether the AnglE Loss turns out to be beneficial for the ordinal and binary task with additional test data. It is also unclear why the AnglE Loss performs better with the cosine similarity than with the angle difference in complex space at test time although the latter is optimized during training.
@@ Appendix B CoMeDi Data
Appendix C Thresholds
Find the thresholds for mapping cosine similarity to ordinal labels for AnglE-WiC+CoMeDi and XL-LEXEME in Table 5. These were induced on the dev data as described in Section 4.4.
Dataset Language AnglE-WiC+CoMeDi XL-LEXEME
CoMeDi ZH .577 .677 .793 .495 .650 .655
CoMeDi EN .325 .483 .612 .418 .607 .682
CoMeDi DE .330 .465 .600 .339 .565 .651
CoMeDi NO .210 .339 .488 .390 .414 .522
CoMeDi RU .255 .491 .615 .249 .511 .749
CoMeDi ES .297 .521 .628 .212 .455 .788
CoMeDi SV .290 .452 .564 .419 .646 .672
MCL-WiC AR .634 .741
MCL-WiC ZH .766 .814
MCL-WiC EN .668 .705
MCL-WiC FR .623 .667
MCL-WiC RU .594 .775
WiC EN .551 .752
Table 5: Thresholds for AnglE-WiC+CoMeDi and XL-LEXEME.
Appendix D Spearman Results
Figure 3 shows model performances measured with Spearman correlation and no thresholding.
Figure 3: Model evaluation with Spearman’s $\rho$ on binary and ordinal test data. CL = Contrastive Loss, COS = Cosine Similarity Loss, CSENT = CoSENT Loss, AnglE = AnglE Loss, XL-LXM = XL-LEXEME.